VIRMOS MMU installed at Paranal

VIRMOS MMU installed at Paranal !!!

The VIRMOS MMU after testing and optimization in Milan and the PAE (Preliminary Acceptance Europe) in April 2000 has been disassembled on July 6, packed, collected by the transporter on July 13 and successfully delivered to the VLT site in Paranal during July-August 2000, to support operation of FORS2.

here is a quick visual journal of the installation ...
(full size pictures can be viewed clicking on table cells)

[JPEG: lab with wall removed] [JPEG: inside of empty lab]
23 July : Arrival of IFCTR team (G.Conti & E.Santambrogio) at Paranal.
#2 The MMU Laboratory with its front wall was removed to allow entrance of the optical table of the laser machine (2.5 tons and around 1.8 x 1.8 meter surface).
#1 Inside of the MMU laboratory at arrival.
[JPEG: lab outside front view] [JPEG: lab outside side view]
#4 View of entire building with main entrance and front wall removed. #3 External area with air conditioning system and prepared to receive the chiller, pressure duplicator and exhauster.
[JPEG: empty lab inside] [JPEG: wall opening from inside]
#6 Another view of the empty room at arrival. #5 The opening in the wall prepared with the platform for transport inside of the MMU Laboratory.
status= in progress
[JPEG: unloading crates] [JPEG: moving crates in]
24 July : Arrival of the MMU crates around noon.
#7 Unloading one of the 10 crates from the lorry (shared with the M3 mirror for UT4). All crates arrived safely : the tilt and shock detectors attached externally to the boxes showed no irregularities.
#9 Moving the crate on the platform for the delicate operation of transport inside of the MMU Laboratory.
[JPEG: unpacking laser] [JPEG: laser unpacked]
#10 The optical bench of the laser machine unpacked, after opening of the sealing bag #13 The optical bench of the laser machine (without laser head) once safely positioned inside the room.
[JPEG: closing wall] [JPEG: laser covered]
#14 Start of wall reconstruction...
successfully ended on the next day
25 July : while completing electrical and LAN connections ...
#15 The control unit of the LPKF machine and the laser head moved inside the room (in background behind the main body of the laser machine protected by the sealing bag).
[JPEG: outside devices] [JPEG: equipment in room]
26 July : works outside the laboratory while the wall is being finished and painted ... OK
#29 The original (provisional) placement (from left under the window to the right) of the pressure duplicator, the exhauster and the chiller. See also picture #28
27 July : first testing of compressed air devices while moving other equipment inside (IC Robot)
28 July : moving other equipment inside (SC and PCs)
#30 view of the equipment [see description below]
[JPEG:equipment inside] [JPEG: Robot view]
#23 The equipment placed inside. The IC Robot with IC holder (front left) and the computers on the back. The Storage Cabinet (front right) with the laser machine and its rack (still with laser head disassembled) on its back. #24 close view of the IC Robot with IC holder and the MHCU computer. Part of the SC protruding from the right.
[JPEG: laser view] [JPEG: Robot front view]
#25 In the foreground the laser machine and its rack (still with laser head disassembled) on the back. On the left on the background the MMCU computer. On the right note the filters (black) and dryer (red ) fixed on the wall near to the window. #27 Front view of the IC Robot with MHCU computer on its left.
PCs and IC Robot status=OK
[JPEG: SC installed] [JPEG: Roughness meter installed]
29 July : continuing the installation ...
#31 Full view of the Storage Cabinet in the front, with bar code cable pivot installed on ceiling.
SC status=OK
30 July : filling of the chiller and testing of outside equipment
31 July : roughness meter setup and software checkout
#94 The roughness meter (on the corner table) and the inspection microscope (on the front desk)
Roughness meter and MHS s/w status=OK
[JPEG: laser head installation] [JPEG: and cleaning]
1 August : arrival of LPKF staff and laser head installation
#47 The laser head (with protection removed) is installed on the optical bench by A.Müller of LPKF
#49 cleaning the granite stand preparing for first switch on.
[JPEG: works at laser head] [JPEG: works at laser head]
2 August : LPKF maintenance protocol and training
#50 ESO staff being trained to replace the pumping Xenon lamp
#51 continuation of the checkout and alignment of the optical system.
[JPEG: end laser installation] [JPEG: laser machine cutting]
#53 Removal of the invar holder frame prior to checkout of air pads and spindles of the motion system #54 Cutting the first test samples at Paranal. The Y-limit switch develops a failure ...
...which is fixed the next day redoing the wiring in the exhauster arm.
[JPEG: outside devices] [JPEG: mask cutting]
3 August : continued work on LPKF laser machine and rearrangement of external equipment
#55 The pressure duplicator has been moved to its final position under the exhauster. Also the pressure of the main air supply has been reduced. status=OK
#58 First VIRMOS mask being manufactured at Paranal.
Laser machine status=OK
LPKF protocol signed by G.Avila of ESO
[JPEG: lab inside view] [JPEG: air filters]
4 August : LPKF training course to ESO staff
#59 the MHCU and MMCU computers, with the laser machine in foreground
#92 adjustment of the input pressure into the duplicator necessary (from 10 to 8 bar) and addition of another filter (6 August) (last unit shown)
[JPEG: outside device housing] [JPEG: outside device front view]
5 August : start construction of housing for external equipment and first meeting with FORS2 team
#91 the material for the housing is being assembled During construction operation on the laser machine is suspended.
#97 The two parts of the housing frame, with (left) exhauster (top) and pressure duplicator (bottom) and (right) chiller
[JPEG: housing frame completed] [JPEG: FORS mask]
6-7 August : housing frame for external equipment done
#99 Housing frame finished. External panels and roof to be mounted later. Laser machine can be switched on.
Housing status=to be completed
8-10 August : definition of MMU operation cycle for FORS2 and training of ESO staff for routine operations
#93 First FORS2 mask cut at Paranal from a Gerber file transmitted by the relevant workstation. Note that FORS2 mask production, unlike VIRMOS which is controlled by MHS, is a manual procedure.
MHS will not be used until VIRMOS arrives at Paranal.
[JPEG: FORS mask container] [JPEG: Robot motor card open]
#95 The bin used to carry FORS2 mask to the telescope for insertion in their Mask eXchange Unit. Note that, unlike VIRMOS where this operation is done with the IC robot, for FORS2 this operation is done at the telescope with no MMU involvement.
The SC and IC Robot will not be used until VIRMOS arrives at Paranal.
11 August : MMU demo to ESO and delivery of Safety Compliance Report
#101 During a test of the IC Robot a failure occurs. Analysis reveals the failure is somewhere on the controller card (shown), which is removed (to be brought to Milano for a revision by Antil).
The failure has been successfully repaired replacing the power supply during a visit by G.Conti in July 2001.
12-14 August : support to FORS2 operations
Masks are manufactured by ESO staff with IFCTR assistance upon reception of Orders from FORS2. IFCTR team leaves to return to Italy. Further FORS2 masks are manufactured on the next day by ESO staff alone.

Summary of MMU status at end of the Preliminary Acceptance Chile

updated after visit in July 2001

MMU room, climatization, power and LAN connections OK
Laser machine OK
MMCU computer and s/w OK
Storage Cabinet OK
IC robot controller card failed repaired July 2001
(power supply replaced)
MHCU computer and s/w OK
roughness meter OK
spare computer and s/w OK
external housing panel walls and roof missing completed later care of ESO
chiller, exhauster and pressure duplicator OK


Back to MMU home